Please help me with this equation: You have a total of $5000 invested in two simple interest accounts. On one account, the annual simple interest rate is 6%. On the second account, the annual simple interest rate is 8%. The total annual interest earned in one yar is $370. How much is invested in each account?
Answer by Theo(13342)

account 1 is 6% per year.
account 2 is 8% per year.
total annual interest in one year is $370.
how much is invested in each account?
let x = amount invested in account 1
let y = amount invested in account 2
total invested = 5000
equations you get are:
x + y = 5000 (equation 1)
.06x + .08y = 370 (equation 2)
solve these 2 equations simultaneously to find your answer.
multiply both sides of first equation by .06 to get:
.06x + .06y = .06*5000
this becomes:
.06x + .06y = 300 (equation 3)
subtract equation 3 from equation 2 to get:
.06x + .08y = 370 (equation 2) minus:
.06x + .06y = 300 (equation 3) equals:
.08y - .06y = 370 - 300 which becomes:
.02y = 70
divide both sides of this equation by 2 to get:
y = 70/.02 = 3500
since x + y = 5000 then x must equal 1500 because 1500 + 3500 = 5000
your answer is:
amount invested at 6% is 1500 and amount invested at 8% is 3500.
total interest is .06 * 1500 + .08 * 3500 = 370.
answer is confirmed as good.
you invest 1500 at 6% and 3500 at 8%